OpenStudy (anonymous):

write the following using only one trig function: cos20cos300+sin20sin300

OpenStudy (mertsj):

\[\cos (a-b)=\cos a \cos b+\sin a \sin b\]

OpenStudy (anonymous):

how we are gonna plug it in?

OpenStudy (mertsj):

Maybe you could just match it up

OpenStudy (anonymous):

how can we find cos20 in unit circle?

OpenStudy (mertsj):

Can you see that a = 20 and b = 300?

OpenStudy (mertsj):

Replace every a with 20 and every b with 300